40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5151     Entry time: Tue Aug 9 03:05:05 2011
Author: kiwamu 
Type: Summary 
Category: General 
Subject: weekly report 

Summary of the week ending Aug 8th.  Number of elog entries = 56


- VAC
 + The vent started Wednesday morning
 + Repositioning of the green periscopes and associated mirrors are done.
 + Got both of the green beams coming out from  the chambers
 + Moved the ETMX suspension tower by -8.09 inch (away from vertex)
 + Fixed the alignment of the ETMX CCD mirrors
 + Recovered the X green beam axis for the latest ETMX position

- SUS
 + oplev centered prior to the vent

- LSC
 + ETMY_TRANS_QPD didn't respond at all, needs to be fixed
 + Old MZ PD (InGaAs 2mm, @29.5MHz) has been modified for REFL33.
  The 11MHz notch circuit is at the amp side instead of the diode side. This is ready for the installation
 + REFL165 PD has been made from the old 166MHz PD.

- IOO
 + IPPOS has been sick since 19th of July, 2011
 + IPANG is clipped on a pick-up mirror on the ETMY table. QPD itself is healthy.
 + The spot positions on the MC mirrors were measured prior to the vent.
   The results are almost the same as before within a few percent difference expect for the MC2 yaw.
 + An attenuator, consisting of two HWPs and a PBS, has been installed on the PSL table for the MC low power state.
 + a 10% BS in front of the MCREFL_RFPD was replaced by a perfect reflector for the low power mode.
 + The incident power for MC was decreased to 20 mW
 + The beam axis going to MC was misalgned due to the attenutor.
    Then the beam was aligned by touching two steering mirrors on the PSL table
 + MC is able to be locked in air. The reflection DC goes from 1.4 to 0.13 V when the MC is locked.

- ABSL
 + With the mass-kicking technique, the arm lengths were measured.
   Xarm =  37.5918 m, Yarm = 37.5425 m.

- Green locking
 + Y green beam is aligned to the Y arm
 + Locking of the Y green is not robust, it needs to be revisited

- OAF
 + Wiener Filtering was applied on the data collected from the X-arm for a duration of 1500 seconds.

- Misc.
 + The hazardous waste people are moving chemicals around outside our door, and have roped off our regular front door.
 + The horizontal trolley drive of the east end crane stopped working. It will be fixed.

ELOG V3.1.3-